# Object Proxy Patterns for Accelerating Distributed Applications

### v1

@@ Abstract
Workflow and serverless frameworks have empowered new approaches to distributed application design by abstracting compute resources. However, their typically limited or one-size-fits-all support for advanced data flow patterns leaves optimization to the application programmer—optimization that becomes more difficult as data become larger. The transparent object proxy, which provides wide-area references that can resolve to data regardless of location, has been demonstrated as an effective low-level building block in such situations. Here we propose three high-level proxy-based programming patterns—distributed futures, streaming, and ownership—that make the power of the proxy pattern usable for more complex and dynamic distributed program structures. We motivate these patterns via careful review of application requirements and describe implementations of each pattern. We evaluate our implementations through a suite of benchmarks and by applying them in three substantial scientific applications, in which we demonstrate substantial improvements in runtime, throughput, and memory usage.
@@ I Introduction
Many researchers have investigated alternative mechanisms for distributed and wide-area data management that avoid this need for shared storage. For example, Linda’s tuple space model provides unified access to a shared distributed memory space [10], DataSpaces provides a similar model for large-scale applications [11, 12], and peer-to-peer systems like the InterPlanetary File System provide decentralized content-addressed file sharing [13]. Another approach to simplifying data sharing is the object proxy paradigm, which provides transparent access and management for shared objects in distributed settings. This mechanism, long used with Java’s Remote Method Invocation (RMI) [14], is also supported in Python via the ProxyStore system [15]. ProxyStore’s transparent object proxies provide lightweight, wide-area references to objects in arbitrary data stores—references that can be communicated cheaply and resolved just-in-time via performant bulk transfer methods in a manner that is transparent to the consumer code. Recent work has shown how by decoupling data flow complexities from control flow-optimized execution engines [16, 17, 18, 19], the object proxy paradigm can simplify implementations of dynamic application structures such as ML model training and inference.
Figure 1: Overview of the three proxy-based data-flow patterns we design.
@@
Each pattern simplifies building sophisticated task-based applications that are to execute across distributed or remote compute resources (e.g., using FaaS platforms or workflow systems). For each, we discuss its requirements and the protocols used to support it and provide an open-source reference implementation at https://github.com/proxystore/proxystore. We choose Python for its popularity and pervasiveness in task-based distributed applications and because it allows us to leverage the proxy model provided by ProxyStore [15]. We evaluate our reference implementation for each pattern on (1) a suite of synthetic benchmarks across various FaaS and workflow systems and (2) our motivating applications, for which we reduce workflow makespan by 36% in 1000 Genomes, improve inference latency by 32% in DeepDriveMD, and optimize proxy lifetime management in MOF Generation.
The rest of this paper is as follows: Sec II introduces our motivating scientific applications; Sec III gives background on the object proxy paradigm; Sec IV outlines the design and implementation of each pattern; Sec V demonstrates synthetic evaluations; Sec VI presents our experiences applying these patterns to our motivating applications; Sec VII provides context about related work; and Sec VIII summarizes our contributions and future directions.
@@ II Motivating Applications
1000 Genomes: This bioinformatics pipeline [20] identifies mutational overlaps within the $2504$ human genomes sequenced by the 1000 Genomes Project [21, 22]. It comprises five stages: (1) fetch files, each containing all Single Nucleotide Polymorphisms (SNPs) in a chromosome, chunk, and process them in parallel to extract SNP variants by individual; (2) merge individuals’ results of the prior stage; (3) score and select SNP variants based on their phenotypic effect; (4) compute overlap of selected SNP variants among pairs of individuals and by chromosome; and (5) compute frequency of overlapping variants. Executing scientific workflows in a FaaS setting may be preferred when access to specialized hardware, such as AI or quantum accelerators, or the ability to rapidly scale up or down is required, but workflow execution on a FaaS system poses challenges because FaaS systems rely on control flow to determine when to submit tasks. From the application perspective, however, the availability of data—the data flow—is the condition upon which tasks can be submitted. We use the 1000 Genomes workflow as an example of the challenges that arise when executing data flow oriented applications on control flow-optimized systems.
DeepDriveMD: Molecular dynamics (MD) simulation acts as a computational microscope [23] to enable the study of complex biomolecular systems [24, 25, 26, 27]. However, many important phenomena are difficult to sample using conventional MD, even with powerful supercomputers [28]. DeepDriveMD [29, 30] implements an emerging HPC paradigm in which machine learning (ML) methods are used to track a simulated state space and guide simulations toward a sampling objective. The DeepDriveMD client submits discrete training, inference, and simulation tasks and receives their results. This pattern causes two challenges. First, all data must flow through the client which limits performance at scale (e.g., data volume or task frequency), so a mechanism is needed to alleviate data flow burdens from the client when possible. Second, repeated tasks perform redundant work. For example, each inference tasks loads the latest ML model from disk, infers using the input batch, and compiles the results which will later become the input to a simulation task. This is inefficient because the same model is loaded multiple times across tasks, tasks may execute on different workers negating cache benefits, and every task incurs non-trivial overheads for scheduling and execution.
@@ III Background
ProxyStore implements this proxy paradigm which we use as the basis for our patterns’ reference implementations. ProxyStore defines the factory, connector, and store constructs. The factory is a callable object that returns the target object when invoked. ProxyStore creates a unique factory for each target object containing the metadata and logic necessary to retrieve the target from a remote location. This factory is used to initialize a proxy, and a proxy is resolved once it has invoked its factory to retrieve and cache the target locally.
@@
The high-level store interface, initialized with a connector, is used to create proxies of objects. A proxy p can be created from a target t by calling Store.proxy(t). This method (1) serializes t, (2) puts the serialized t in the mediated channel via the connector, (3) creates a factory with the appropriate metadata about t and the store/connector used, (4) initializes a proxy with the factory, and (5) returns the proxy.
@@ IV Proxy Patterns
We describe the design of each of the three advanced programming patterns that build on the aforementioned distributed object proxy base. We discuss the details of our reference implementation which we choose to build on ProxyStore, and Fig 2 describes how these patterns fit into the existing ProxyStore stack. These patterns are not mutually exclusive, but we discuss each in isolation for clarity.
@@ IV-A Distributed Futures
1from proxystore.connectors import FooConnector
2from proxystore.store import Store
3from proxystore.store.future import Future
4
5with Store('example', FooConnector()) as store:
6 future: Future[str] = store.future()
7 proxy = future.proxy()
8
9 future.set_result('value')
10 assert proxy == 'value'
Listing 1: Example usage of the ProxyFuture interface. A proxy created from a Future instance will block implicitly on the result of the future when needed. This interface abstracts the low-level communication away from the processes which set the result or consume the future.
We implement this behavior by extending ProxyStore’s Store interface to expose a future() method that returns a Future object. The Future class exposes two main methods: set_result(obj: T), which sets the result of the future to an object of type T and proxy(), which returns a Proxy[T]. When a proxy created via Future.proxy() is resolved, the proxy blocks until the target value has been set via a call to Future.set_result(), as shown in 1. Internally, communication between a Future and any child proxy(s) is handled via the Store instance used to create the Future. Thus, a future and associated proxies can be serialized and sent to arbitrary processes on arbitrary machines. (In contrast, many standard-library future implementations use non-serializable async, thread, and inter-process synchronization mechanisms, while RPC-based futures are only resolvable within the RPC framework.) The self-contained properties of the proxy mean that all logic for communication and resolution are embedded within the future and proxy; the creator of the future chooses the communication method on behalf of the process(es) which might set or consume the result of the future.
@@ IV-B Object Streaming
High-performance stream processing applications dispatch remote compute tasks on objects consumed from a stream, but task dispatch can quickly become a bottleneck with high throughput streams, as in [29, 30, 35]. Consider the application in Fig 4, where process $A$ is a data generator that streams chunks of data (i.e., arbitrary Python objects) to a dispatcher process $B$, which for each data chunk dispatches a compute task on a remote process $C$. Note that while the dispatcher consumes from the stream, it does not need the actual chunk of data; rather, it only needs to know that a chunk is ready (and potentially have access to user-provided metadata) in order to dispatch the task that will actually consume the chunk. We design a system called ProxyStream to enable scalable applications of this pattern. At its core, ProxyStream uses a stream of proxies, rather than data chunks. Bulk data are only transmitted between the data generator and the process/node computing on the proxy of the chunk, bypassing the intermediate dispatching process. ProxyStream optimizes for both metadata and bulk data transfer, has broad execution engine compatibility, provides a self-describing data format, and supports various communication modules to take advantage of high-performance networking stacks.
1from proxystore.connectors import FooConnector
2from proxystore.store import Store
3from proxystore.stream import StreamProducer
4from proxystore.stream import KafkaPublisher
5
6store = {'topic': Store(..., FooConnector(...))}
7publisher = KafkaPublisher(...)
@@
9with StreamProducer(publisher, store) as producer:
@@
Listing 2: Example stream producer code using Kafka for metadata and an arbitrary FooConnector for bulk data transfer.
1from proxystore.proxy import Proxy
2from proxystore.stream import StreamConsumer
3from proxystore.stream import KafkaSubscriber
4
5subscriber = KafkaSubscriber('topic', ...)
6
7with StreamConsumer(subscriber) as consumer:
@@
Listing 3: Example stream consumer code using Kafka for metadata. The bulk data transfer mechanism is determined by the producer (a FooConnector in 2) and encoded in the event metadata.
ProxyStream provides two high-level constructs, the StreamProducer and StreamConsumer, that combine a message stream broker for low-latency event metadata propagation and a mediated communication channel for efficient bulk data transfer. A StreamProducer is initialized with a Publisher and a ProxyStore Store (2). The Publisher defines a protocol for sending event messages to a stream. We provide shims to many popular event streaming systems (Kafka [44], Redis Pub/Sub and Queues [38], ZeroMQ [45]) which implement the Publisher protocol. When a new object and optional metadata are sent to the StreamProducer, (1) the object is put in the store, (2) a new event containing the user provided metadata and information about where the object is stored is created, and (3) the event is published via the Publisher.
@@ IV-C Ownership
Listing 4: Proxy ownership model interfaces and functions. Functions are preferred over methods on the associated proxy reference types to prevent unintentionally clobbering a method of the same name on the target object.
@@
Execution engines typically use futures to encapsulate the asynchronous execution of a task. Thus, we use callbacks on the task result futures to indicate that the references associated with a task have gone out of scope. The primary limitation of this approach is that each execution engine has a different syntax for submitting a task and getting back a future. Rather than modify the engines, we provide a set of shims that appropriately parse task inputs and construct a callback on the task’s future that will propagate the necessary information about references going out of scope. We believe this approach to be satisfactory for this proof-of-concept implementation because it gives flexibility to use this system with a wide range of compute frameworks, but further features could be supported if execution engines closely integrate this model. The ownership model is not fault-tolerant when the client crashes in a manner which prevents garbage collection, but the model is compatible with fault-tolerant execution engines such as those that automatically rerun tasks on failure.
So far, we have constricted ourselves to tasks (i.e., function invocations) as the only region of code over which we can define a lifetime; thus, all references to an object are equal to the lifetime of the single task invoked on that reference. Yet a workflow application may employ more complex lifetimes. For example, a lifetime could be assigned to a set of tasks that are a subgraph of the global DAG, and a programmer might want to define references to global objects that are associated with this custom lifetime. We provide the Lifetime construct, an alternative to proxy references, for managing object lifetimes in more complex scenarios. A lifetime, attached to a proxy when created, will clean up associated objects once the lifetime has ended. We provide context-manager, time-leased, and static Lifetime implementations.

### v2

@@ Abstract
Workflow and serverless frameworks have empowered new approaches to distributed application design by abstracting compute resources. However, their typically limited or one-size-fits-all support for advanced data flow patterns leaves optimization to the application programmer—optimization that becomes more difficult as data become larger. The transparent object proxy, which provides wide-area references that can resolve to data regardless of location, has been demonstrated as an effective low-level building block in such situations. Here we propose three high-level proxy-based programming patterns—distributed futures, streaming, and ownership—that make the power of the proxy pattern usable for more complex and dynamic distributed program structures. We motivate these patterns via careful review of application requirements and describe implementations of each pattern. We evaluate our implementations through a suite of benchmarks and by applying them in three meaningful scientific applications, in which we demonstrate substantial improvements in runtime, throughput, and memory usage.
@@ I Introduction
Many researchers have investigated alternative mechanisms for distributed and wide-area data management that circumvent these limitations of shared storage. For example, Linda’s tuple space model provides unified access to a shared distributed memory space [10], DataSpaces provides a similar model for large-scale applications [11, 12], and peer-to-peer systems like the InterPlanetary File System provide decentralized content-addressed file sharing [13]. Another approach to simplifying data sharing is the object proxy paradigm, which provides transparent access and management for shared objects in distributed settings. This mechanism, long used with Java’s Remote Method Invocation (RMI) [14], is also supported in Python via the ProxyStore system [15]. ProxyStore’s transparent object proxies provide lightweight, wide-area references to objects in arbitrary data stores—references that can be communicated cheaply and resolved just-in-time via performant bulk transfer methods in a manner that is transparent to the consumer code. Recent work has shown how by decoupling data flow complexities from control flow-optimized execution engines [16, 17, 18, 19, 20], the object proxy paradigm can simplify implementations of dynamic application structures such as ML model training and inference.
Figure 1: Overview of the three proxy-based data flow patterns we design.
@@
Each pattern simplifies building sophisticated task-based applications that are to execute across distributed or remote compute resources (e.g., using FaaS or workflow systems). For each, we discuss its requirements and the protocols used to support it. Our reference implementations extend ProxyStore [15], our prior work, to leverage the existing low-level proxy model within Python, a popular and pervasive language for task-based distributed applications. The implementations are available within ProxyStore v0.6.5 and later, available on GitHub [21] and PyPI [22]. We evaluate our reference implementation for each pattern using (1) synthetic benchmarks across various FaaS and workflow systems and (2) our motivating applications, for which we reduce workflow makespan by 36% in 1000 Genomes, improve inference latency by 32% in DeepDriveMD, and optimize proxy lifetimes in MOF Generation.
The rest of this paper is as follows: Sec II introduces our motivating scientific applications; Sec III gives background on our prior work, ProxyStore; Sec IV outlines the design and implementation of each pattern; Sec V demonstrates synthetic evaluations; Sec VI presents our experiences applying these patterns to our motivating applications; Sec VII provides context about related work; and Sec VIII summarizes our contributions and future directions.
@@ II Motivating Applications
1000 Genomes: This bioinformatics pipeline [23] identifies mutational overlaps within the $2504$ human genomes sequenced by the 1000 Genomes Project [24]. It comprises five stages: (1) fetch files, each containing all Single Nucleotide Polymorphisms (SNPs) in a chromosome, chunk, and process them in parallel to extract SNP variants by individual; (2) merge individuals’ results of the prior stage; (3) score and select SNP variants based on their phenotypic effect; (4) compute overlap of selected SNP variants among pairs of individuals and by chromosome; and (5) compute frequency of overlapping variants. Executing scientific workflows in a FaaS setting may be preferred when access to specialized hardware, such as AI or quantum accelerators, or the ability to rapidly scale up or down is required, but workflow execution on a FaaS system poses challenges because FaaS systems rely on control flow to determine when to submit tasks. From the application perspective, however, the availability of data—the data flow—is the condition upon which tasks can be submitted. We use the 1000 Genomes workflow as an example of the challenges that arise when executing data flow oriented applications on control flow-optimized systems.
DeepDriveMD: Molecular dynamics (MD) simulation acts as a computational microscope [25] to enable the study of complex biomolecular systems. However, many important phenomena are difficult to sample using conventional MD, even with powerful supercomputers [26]. DeepDriveMD [27, 28] implements an emerging HPC paradigm in which machine learning (ML) methods are used to track a simulated state space and guide simulations toward a sampling objective. The DeepDriveMD client submits discrete training, inference, and simulation tasks and receives their results. This pattern causes two challenges. First, all data must flow through the client which limits performance at scale (e.g., data volume or task frequency), so a mechanism is needed to alleviate data flow burdens from the client when possible. Second, repeated tasks perform redundant work. For example, each inference tasks loads the latest ML model from disk, infers using the input batch, and compiles the results which will later become the input to a simulation task. This is inefficient because the same model is loaded multiple times across tasks, tasks may execute on different workers negating cache benefits, and every task incurs non-trivial overheads for scheduling and execution.
@@ III ProxyStore
ProxyStore [15] implements this proxy paradigm which we use as the basis for our patterns’ reference implementations. ProxyStore defines the factory, connector, and store constructs. The factory is a callable object that returns the target object when invoked. ProxyStore creates a unique factory for each target object containing the metadata and logic necessary to retrieve the target from a remote location. This factory is used to initialize a proxy, and a proxy is resolved once it has invoked its factory to retrieve and cache the target locally.
@@
The high-level store interface, initialized with a connector, is used to create proxies of objects. A proxy p can be created from a target t by calling Store.proxy(t). This method (1) serializes t using the default ProxyStore or user-provided serializer [42], (2) puts the serialized t in the mediated channel via the connector, (3) creates a factory with the appropriate metadata about t and the store/connector used, (4) initializes a proxy with the factory, and (5) returns the proxy. This process incurs some overhead but is trivial for larger objects. Prior work [15, 16] found the performance benefits of proxies to outweigh proxy creation and resolution overhead for objects larger than $\sim$10 kB; the exact threshold depends on many factors (e.g., connector choice, execution engine).
@@ IV Proxy Patterns
We describe the design of each of the three advanced programming patterns that build on the aforementioned distributed object proxy base. We discuss the details of our reference implementations that extend ProxyStore, and Fig 2 describes how these patterns fit into the existing ProxyStore stack. These patterns are not mutually exclusive, but we discuss each in isolation for clarity.
@@ IV-A Distributed Futures
1from dask.distributed import Client
2from proxystore.connectors.foo import FooConnector
3from proxystore.store import Store
4from proxystore.store.future import Future
5
6def producer(future: Future[str]) -> None:
7 future.set_result('value')
8
9def consumer(data: Proxy[str]) -> None:
10 assert data is 'value'
11
12with Store('example', FooConnector()) as store:
13 client = Client(...)
14 future: Future[str] = store.future()
15
16 t1 = client.submit(producer, future)
17 t2 = client.submit(consumer, future.proxy())
18
19 t1.result(), t2.result()
Listing 1: Example usage of the ProxyFuture interface within tasks executed by Dask. A proxy created from a Future will block implicitly on the result of the future when needed. This interface abstracts the low-level communication away from the functions which set the result or consume the proxy.
@@
We implement this behavior by extending ProxyStore’s Store interface to expose a future() method that returns a Future object. The Future class exposes two main methods: set_result(obj: T), which sets the result of the future to an object of type T and proxy(), which returns a Proxy[T]. When a proxy created via Future.proxy() is resolved, the proxy blocks until the target value has been set via a call to Future.set_result(), as shown in 1. Use of ProxyFutures does not affect when a successor task starts; scheduling is still managed by the execution engine and/or user application. ProxyFutures are best integrated at the application level so that developers can optimize task execution per their application requirements and to express more complex data dependencies than typically supported by execution engines.
Internally, communication between a Future and any child proxy(s) is handled via the Store used to create the Future. Thus, a future and associated proxies can be serialized and sent to arbitrary processes on arbitrary machines. In contrast, many standard-library future implementations use non-serializable async, thread, and inter-process synchronization mechanisms (e.g., std::future in C++ [43], concurrent and async futures in Python [44]), while RPC-based futures are only resolvable within the RPC framework (Dask futures [1] or Ray ObjectRefs [45]). The self-contained properties of the proxy mean that all logic for communication and resolution are embedded within the future and proxy; the future creator chooses communication methods on behalf of the process(es) which might set or consume the result of the future.
@@ IV-B Object Streaming
1from globus_compute_sdk import Executor
2from proxystore.stream import (StreamProducer,
3 StreamConsumer, KafkaPublisher, KafkaSubscriber)
4
5def producer() -> None:
6 store = {'topic': Store(..., FooConnector(...))}
7 publisher = KafkaPublisher(...)
@@ 3 StreamConsumer, KafkaPublisher, KafkaSubscriber)
9 with StreamProducer(publisher, store) as producer:
@@
12
13def consumer() -> None:
14 subscriber = KafkaSubscriber('topic', ...)
15
16 with StreamConsumer(subscriber) as consumer:
@@
19
20with Executor('<Endpoint UUID>') as client:
21 t1 = client.submit(producer)
22 t2 = client.submit(consumer)
23
24 t1.result(), t2.result()
Listing 2: Example using the ProxyStream interfaces to stream data between two tasks executed remotely using Globus Compute. A Kafka broker is used for metadata and an arbitrary FooConnector for bulk data transfer.
High-performance stream processing applications dispatch remote compute tasks on objects consumed from a stream, but task dispatch can quickly become a bottleneck with high throughput streams [27, 28, 33]. Consider the application in Fig 4, where process $A$ is a data generator that streams chunks of data (i.e., arbitrary Python objects) to a dispatcher process $B$, which for each data chunk dispatches a compute task on a remote process $C$. Note that while the dispatcher consumes from the stream, it does not need the actual chunk of data; rather, it only needs to know that a chunk is ready (and potentially have access to user-provided metadata) in order to dispatch the task that will actually consume the chunk. We design a system called ProxyStream to enable scalable applications of this pattern. At its core, ProxyStream uses a stream of proxies, rather than data chunks. Bulk data are only transmitted between the data generator and the process/node computing on the proxy of the chunk, bypassing the intermediate dispatching process. ProxyStream optimizes for both metadata and bulk data transfer, has broad execution engine compatibility, provides a self-describing data format, and supports various communication modules to take advantage of high-performance networking stacks.
ProxyStream provides two high-level constructs, the StreamProducer and StreamConsumer, that combine a message stream broker for low-latency event metadata propagation and a mediated communication channel for efficient bulk data transfer. A StreamProducer is initialized with a Publisher and a ProxyStore Store. The Publisher defines a protocol for sending event messages to a stream. We provide shims to many popular event streaming systems (Kafka [46], Redis Pub/Sub and Queues [36], ZeroMQ [47]) which implement the Publisher protocol. When a new object and optional metadata are sent to the StreamProducer, (1) the object is put in the store, (2) a new event containing the user provided metadata and information about where the object is stored is created, and (3) the event is published via the Publisher.
@@
ProxyStream can be integrated at the application or framework level. 2 depicts use of ProxyStream within a Globus Compute application; we integrate ProxyStream within the DeepDriveMD framework for the evaluation in Sec VI.
@@ IV-C Ownership
Listing 3: Proxy ownership model interfaces and functions. Functions are preferred over methods on the associated proxy reference types to prevent unintentionally clobbering a method of the same name on the target object.
@@
Execution engines typically use futures to encapsulate the asynchronous execution of a task. Thus, we use callbacks on the task result futures to indicate that the references associated with a task have gone out of scope. The primary limitation of this approach is that each execution engine has a different syntax for submitting a task and getting back a future. Rather than modify each engine, we provide a set of shims that appropriately parse task inputs and construct a callback on the task’s future that will propagate the necessary information about references going out of scope. The StoreExecutor, an interface provided by ProxyStore, wraps an execution engine client (e.g., a Globus Compute, Dask, or Parsl client) and automatically proxies task parameters and results based on user-defined policies and manages references associated with tasks [42]. The StoreExecutor is easy to use, but applications requiring more fine-grain control can use the API in 3.
The ownership model is not fault-tolerant when the client crashes in a manner which prevents garbage collection, but the model is compatible with fault-tolerant execution engines such as those that automatically rerun tasks on failure. Since only a single RefMutProxy can exist, the ownership model is not optimal for applications with many concurrent writers to the same object; a database, for example, may be more suitable.
So far, we have constricted ourselves to tasks (i.e., function invocations) as the only region of code over which we can define a lifetime; thus, all references to an object are equal to the lifetime of the single task invoked on that reference. Yet a workflow application may employ more complex lifetimes. For example, a lifetime could be assigned to a set of tasks that are a subgraph of the global DAG, and a programmer might want to define references to global objects that are associated with this custom lifetime. Using proxy references is a valid solution but would require additional code to manage and map references to the scopes contextual to the application.
1from proxystore.connectors import FooConnector
2from proxystore.store import Store
3from proxystore.store.lifetimes import LeaseLifetime
4
5with Store('example', FooConnector()) as store:
6 lease = LeaseLifetime(store, expiry=10)
7 proxy = store.proxy('value', lifetime=lease)
8 lease.extend(5)
9 time.sleep(20)
10 assert lease.done()
11 # Object associated with the proxy has been removed
Listing 4: Example usage of lifetimes when creating a proxy. A Lifetime instance represents a physical or logical scope that will clean up all resources (i.e., objects) that were associated with the lifetime when closed.
We provide the Lifetime construct, an alternative to proxy references, for managing object lifetimes in more complex scenarios. A lifetime, attached to one or more proxies upon proxy creation, will clean up associated objects once the lifetime has ended. We provide three Lifetime types and the API can be extended to implement new types. The context-manager lifetime enables mapping proxy lifetimes to discrete segments of code, the time-leased lifetime will clean up associated objects once the lease has expired and not been extended, and the static lifetime persists objects for the remainder of the program. 4 provides a time-leased lifetime example.